CAVITY-ENHANCED ABSORPTION SPECTROSCOPY WITH A MODE-LOCKED FEMTOSECOND LASER

\underline{TITUS GHERMAN}, S. KASSI, A. CAMPARGUE and D. ROMANINI, Laboratoire de Spectrométrie Physique – CNRS UMR 5588, Université J. Fourier–Grenoble I, B.P. 87 – 38402 Saint Martin d’Hères Cedex, FRANCE.

The wide spectral coverage and high sensitivity of Mode-Locked Cavity Enhanced Absorption Spectroscopy (ML-CEAS) are illustrated by the observation of a very high overtone transition of C_{2}H_{2} in the blue spectral region (420 nm, 8 quanta of CH stretch excitation), which was easily accessed by frequency-doubling a modelocked femtosecond Ti:Sapphire laser. The detection limit is about 10^{−8}/cm. The rotational analysis of this ∑_{u}^{+} − ∑_{g}^{+} parallel band, centred at 23813.244 cm-1, is presented and the vibrational assignment is discussed. T.Gherman, S. Kassi, A.Campargue, D. Romanini Chem. Phys. Lett. 383, 353-358 (2004);